A simple microscope is rated 5 X for a normal relaxed eye. What will be its magnifying power for a relaxed farsighted eye whose near point is 40 cm?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. The simple microscope has magnification of 5 for normal relaxed eye (D = 25 cm). Because, the eye is relaxed the image is formed at infinity (normal adjustment) So, m = 5 = D/f = 25/f ⇒ f = 5 cm For the relaxed farsighted eye, D = 40 cm So, m = D/f = 40/5 = 8 So, its magnifying power is 8X.